OpenStudy (anonymous):

HELP PLEASE Find the equation of the circle given the following information: Center (4, -5) and Radius = 3

OpenStudy (anonymous):

Jlastino you are my hero lol :)

OpenStudy (jlastino):

Given the center of the of the circle (h,k) and the radius, the equation is (x-h)^2 + (y-k)^2 = r^2 so what's the equation? (I'm gonna go to an anime convention so i won't stay much long here haha)

OpenStudy (anonymous):

is 4 and -5 the h and k?

OpenStudy (jlastino):

yeah

OpenStudy (anonymous):

ok so (x-4)^2 +(y-5)=3^2

OpenStudy (jlastino):

hmm there's something wrong with (y-5)^2 k= -5 so y-(-5) would become...?

OpenStudy (anonymous):

positive 5

OpenStudy (jlastino):

so what's the right equation? :D

OpenStudy (anonymous):

(x-4)^2 +(y+5)=3^2

OpenStudy (anonymous):

^2 sorry

OpenStudy (jlastino):

oh ok just simplify 3^2 and that would be the correct answer

OpenStudy (jlastino):

(x-4)^2 + (y+5)^2 =9

OpenStudy (anonymous):

right sorry. um so thats the answer :)

OpenStudy (jlastino):

yes :D

OpenStudy (anonymous):

thank you SO much. Have a good time at your convention. What kind of anime do you like?

OpenStudy (jlastino):

haha this isn't supposed to be a chatbox but some anime I like include : ranma ,evangelion, samurai champloo, steins gate, love hina, another, detective conan. :))

OpenStudy (anonymous):

haha i have never heard of any of those lol

OpenStudy (anonymous):

ive heard of dragon ball z :D

OpenStudy (jlastino):

oh that I only know about goku gohan and vegeta haha i wasn't able to watch much of that anime 'cause i wasn't interested in anime back then...more of cartoons xD

OpenStudy (anonymous):

haha yea my fiance is so into it. anyway, have fun :D